OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -83?

OpenStudy (anonymous):

(-83-13)/16=-6(common difference) a35=a1+34*common difference a35=13+34*(-6) a35=-191

OpenStudy (anonymous):

Thank you, i have just one more, i know i said that, i had four i couldnt get.

OpenStudy (anonymous):

ok. i'm waiting :)

OpenStudy (anonymous):

i put it in a new one so i can give you bet response